OpenStudy (anonymous):

A flask contains 119.99 g of NaOH. In order to make a 4.00 M NaOH solution, water should be added until what volume is reached?

OpenStudy (matt101):

Similar procedure as the one we just did. How would you solve this one?

OpenStudy (thadyoung):

You can also use density of sodium hydroxide to get volume.

OpenStudy (anonymous):

I really have no idea. I'm sorry. I'm pretty stressed out today

OpenStudy (matt101):

No worries (PS. @thadyoung, NaOH is dissolving in the water, so it won't contribute significantly to the volume of solution) For most questions like this, looking at your equations can help you decide how to solve them. The equations we will need are the ones for moles (n) and concentration (C): \[n={m \over MM} \space \space \space \space \space \space \space C={n \over V}\] Our unknown is volume (V), so we know we'll be ending with that second equation and solving for V. We have concentration (C) given, so we need just some way of finding n. That's where the first equation comes in. We have a mass (m) given in the equation, and we can easily figure out the molar mass (MM), so we can use that information to find moles, then use moles to find volume! See if you can do this one now!

OpenStudy (anonymous):

So it'll be 4.0m/39.997? @matt101

OpenStudy (matt101):

Back up - first find moles: mass divided by molar mass gives you...?

OpenStudy (anonymous):

119.99/39.997=2.99

OpenStudy (matt101):

Right. We have 2.99 moles of of NaOH. Now that we have n and C, we can solve for V in that second equation. What do you get for V?

OpenStudy (anonymous):

4.0x2.99=11.96

OpenStudy (matt101):

\[C={n \over V} \rightarrow V = {n \over C}\] Solve for V.

OpenStudy (anonymous):

Since C is a division problem wouldn't you have to do the opposite to find V?

OpenStudy (matt101):

Not if you rearrange the equation so that you're solving for V - see what I did above. To isolate V, multiply both sides of the equation by V to get V out of the denominator: \[V \times C = {n \times V \over V}\] V's reduce out on the right side, leaving: \[V \times C = n\] Now just divide both sides by C to get V on its own: \[{V \times C \over C}= {n \over C}\] \[V={n \over C}\] Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. So it's 4.0/2.99=1.33?

OpenStudy (anonymous):

No that's not it. it's supposed to come out to 750mL

OpenStudy (matt101):

You have it backwards - it should be 2.99/4.0.

OpenStudy (anonymous):

That comes out to .7475 Now what?

OpenStudy (anonymous):

@Somy Can you please help? This is my last problem and I'm running short on time

OpenStudy (matt101):

Yes that's it! if you round to the nearest hundredth you get 0.75. Remember, this number is in litres (L). 0.75 L = 750 mL!

OpenStudy (somy):

its done tho? :)

OpenStudy (anonymous):

That's what I get for my computer freaking out on me and not loading stuff

OpenStudy (anonymous):

Thanks everybody!

OpenStudy (somy):

:D